OpenStudy (anonymous):

find lim (x,y) -> (0,0) of (x+y)(1/x-1/y)

OpenStudy (anonymous):

\[\lim_{x,y \rightarrow (0,0)} (x+y)(\frac{ 1 }{x } + \frac{ 1 }{ y})\]

OpenStudy (amistre64):

x+y 1/x - 1/y x^2y + xy^2 y - x might not be easier, but at least gets rid of the fractions

OpenStudy (anonymous):

so, i simplied to: y^2-x^2/(xy), I then used polar coordinates and got to r/cos(theta)sin(theta)

OpenStudy (amistre64):

the limit of a multivariable only exists if its the same from ALL directions ... which can be an utter pain to assess

OpenStudy (anonymous):

not sure what you did above?

OpenStudy (amistre64):

mulitplied top and bottom by xy

OpenStudy (amistre64):

and your post has -, but the latex has +

OpenStudy (amistre64):

lol, i thought i saw a division bar in the ascii :/

OpenStudy (anonymous):

my bad, its minus

OpenStudy (amistre64):

my idea is to test the limit in key directions, like y=x and such

OpenStudy (anonymous):

ye, thats a good idea.I was trying to pick a direction to contradict eachother.

OpenStudy (amistre64):

\[\lim_{x,y \rightarrow (0,0)} (x+x)(\frac{ 1 }{x } + \frac{ 1 }{ x})=1\]

OpenStudy (anonymous):

its a minus, remember

OpenStudy (anonymous):

lim (x,y) -> (0,0) of (x+y)(1/x-1/y)

OpenStudy (amistre64):

ugh, too early to remember ...

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

maybe y = 2x. in (x^2y + xy^2)(y - x)